OpenStudy (anonymous):

Two trains leave stations 440 miles apart at the same time and travel toward eachother. One train travels 110 miles per hour while the other travels 90 miles per hour. How long will it take for the two trains to meet. don't do any rounding.

OpenStudy (whpalmer4):

In one hour, how many miles do the two trains cover? How many hours does it take to cover 440 miles at that rate?

OpenStudy (anonymous):

Is this a 110+40x=440 type of question..not sure what the math term is, lol forgot.

OpenStudy (whpalmer4):

okay, how far does the fast train go in one hour? how far does the slow train go in one hour? together, how many miles of track have they covered in 1 hour?

OpenStudy (whpalmer4):



OpenStudy (sleepyjess):

(110+90)x=440 200x=440 x=440/200 x=11/5 x=2.2

OpenStudy (whpalmer4):

Yes, that is correct. Of course, my 10-year-old did it without using algebra :-)

OpenStudy (whpalmer4):